OpenStudy (anonymous):

When the quadratic equation y = (x − 5)(2x + 3) is written in standard form, what is the value of the coefficient "b"? 7 −7 2 −15

OpenStudy (jack1):

so, use FOIL technique to expand out the brackets then pick your answer, knowing that y = ax^2 + bx + c

OpenStudy (jack1):

do you understand FOIL technique...?

OpenStudy (anonymous):

To find the answer we can write : \[(x-5)(2x+3)=2x^2+3x-10x-15=2x^2-7x-15\] Then : \[a=2~~b=-7~~~c=-15\]